# A discrete approach to Dirichlet L-functions via spectral models: special values and zeros

### v1

@@ 3 Asymptotics of discrete Dirichlet $L$-functions
$\zeta_{\mathbb{Z}/n\mathbb{Z}}(s)$ $=n\zeta_{\mathbb{Z}}(s)+n^{2s}\left(\sum_{k=0}^{m}a_{k}(s)\zeta_{\mathbb{R}/\mathbb{Z}}(s-k)n^{-2k}+\mathcal{O}(n^{-2-2m})\right).$
@@
The following asymptotic expression is obtained in [F16].
@@
Let $\chi$ be a even, non-principal Dirichlet character of modulus $q>1$. Then, for any fixed $s\in\mathbb{C}\setminus\{\frac{1}{2},\frac{3}{2},\frac{5}{2},\ldots\}$ and any integer $m\geq 0$, the following asymptotic expansion holds as $n\to\infty$:
$L_{n}(s,\chi)=2\left(\frac{2\pi}{qn}\right)^{-2s}\left[\sum_{k=0}^{m}a_{k}(s)\left(\frac{2\pi}{qn}\right)^{2k}L(2(s-k),\chi)+\mathcal{O}(n^{-2-2m})\right],$
@@
$=n^{2s}\left(\sum_{k=0}^{m}a_{k}(s)\left(\frac{1}{-2\pi s}\right)^{a}\sum_{r=1}^{q-1}\chi(r)(\partial_{\theta}^{a}\zeta_{\mathbb{R}/\mathbb{Z}})(r/q,s-k)n^{a-2k}+\mathcal{O}(n^{a-2-2m})\right)$
$=2n^{2s}\left(\sum_{k=0}^{m}a_{k}(s)\left(\frac{(2s-k)}{s}\right)^{a}\left(\frac{q}{2\pi}\right)^{2(s-k)+a}L(a+2(s-k),\chi)n^{a-2k}+\mathcal{O}(n^{a-2-2m})\right)$
$=2\left(\frac{qn}{2\pi}\right)^{2s}\left(a_{k}(s)\left(\frac{2(s-k)}{s}\right)^{a}L(a+2(s-k),\chi)\left(\frac{qn}{2\pi}\right)^{a-2k}+\mathcal{O}(n^{a-2-2m})\right)$
Which concludes the proof of both Theorem 1.1 and Theorem 3.2, recalling that
@@
In fact, in the next section, what will be most important for us is the two first terms of the asymptotic, that is the case $m=1$ in Theorem 1.1. Precisely the asymptotic for $m=1$ is
@@
Let $\chi$ be a non-principal, even and real Dirichlet character of modulus $q>1$. Fix $0<s<1$. Then $L(\chi,s-2)<0$ and it follows from the asymptotics in Theorem 3.2 that if $L_{n}(s,\chi)\geq 0$ for a subsequence of $n\rightarrow\infty$, then $L(\chi,s)>0$ and in particular non-zero.
Let $\chi$ be an odd and real Dirichlet character of modulus $q$. Fix $0<s<1$. The second term in the asymptotics in Theorem 1.1 now has a different sign, since (s3)/6<0(s-3)/6<0. It follows that if $L_{n}(s,\chi)\geq 0$ for a subsequence of $n\rightarrow\infty$, then $L(\chi,s)\geq 0$.
Since Siegel zeros are known to be simple for real characters, they cannot exist if $L(\chi,s)\geq 0$ for all $0<s<1$ (or actually, it is enough with a certain interval $1-\epsilon<s<1$.)
@@
and it satisfy the functional equation
@@
We now specialize the previous theorem when $\chi$ is real, odd and for real zeroes of $L$. Recall that for such a $\chi$, Gauss himself proved that $w_{\chi}=1$. Therefore, as in the previous proof, for real $0<s<1$, if $\xi(s,\chi)\neq 0$, then
@@ 5.1 No real zero for Dirichlet $L$-functions with heat positive character
The Mellin transform of the sum in the hypothesis gives $L_{n}(s,\chi)\cdot\Gamma(s)$. The $\Gamma$ factor is positive on $s>0$, and hence heat positivity implies that $L_{n_{i}}(s,\chi)>0$. Therefore, since $L(s-2,\chi)<0$ for $0<s<1$ as remarked above, the asymptotics (10) force $L(\chi,s)>0$. ∎
@@ 5.2 First examples of heat positive characters
$(\Psi_{n}\rho)(\omega)=4\sin^{2}\left(\frac{\pi\omega}{qn}\right)\rho(\omega)$
. One computes for $\omega\in\mathbb{Z}/qn\mathbb{Z}$,
@@
since we consider $\mathbb{Z}/qn\mathbb{Z}$ as discrete, the dual measure (with respect to Fourier) is therefore the compact one on $\mathbb{Z}/qn\mathbb{Z}$. Taking the trace of this solution viewed as a operator acting on $L^{2}(\mathbb{Z}/qn\mathbb{Z})$ by convolution amounts to putting $x=0$ and multiplying by $qn$, which gives
@@ 5.3 Further examples of heat positive characters
Consider the operator $S\coloneqq 3-\Delta$, defined by $(Sf)(x)\coloneqq f(x-1)+f(x)+f(x+1)$, which is the generator of the random walk on $\mathbb{Z}/N\mathbb{Z}$ allowing the walk to stand still. Let us also $\partial$ be the forward difference operator $(\partial f)(x)\coloneqq f(x+1)-f(x)$. Then $K(t)=e^{-3t}\exp(tS)\delta_{0}$, and therefore we need only to analyze $\phi(t)\coloneqq\exp(tS)\delta_{0}$. But clearly,
@@ 6 Special values
In this section, we prove Theorem 1.2, and by doing so we link special values of both discrete zeta functions and discrete $L$-functions with special values of the Riemann zeta function and Dirichlet $L$-functions, respectively. For this, we first use the Laplace transform of the discrete heat kernel on the $\theta$-bundle over $\mathbb{Z}/n\mathbb{Z}$ together with the characteristic polynomial of the bundle Laplacian $\Delta_{n,\theta}$ (namely, the Chebyshev polynomials) to obtain a polynomial behavior of the discrete special value. Combining this polynomial behavior with the asymptotics of Theorem 2.1, we obtain closed formulas for the zeta functions associated. By using standard formulas for $L$-functions and specializing $\theta$ accordingly, we obtain the desired relation between discrete and continuous zeta and $L$-functions.
@@ 6.1 Formulas
In terms of discrete zeta function on the $\theta$-bundle, Theorem 1.2 says that for any $0<\theta<1$ and integers $n,p\geq 1$, we have
$\zeta_{\mathbb{Z}/n\mathbb{Z}}(p,\theta)=\sum_{k=0}^{p}a_{p-k}(p)\zeta_{\mathbb{R}/\mathbb{Z}}(k,\theta)n^{2k}.$ (23)
As above, the maps $s\mapsto a_{k}(s)$ are polynomial functions of degree $k$, given by the Taylor expansion
@@
For any integers $n,p\geq 1$, the following holds:
$\zeta_{\mathbb{Z}/n\mathbb{Z}}(p)=\sum_{k=0}^{p}a_{p-k}(p)\zeta_{\mathbb{R}/\mathbb{Z}}(k)n^{2k}.$
@@ 6.3 Laplace transform of the bundle heat kernel
$\sum_{j=0}^{n-1}\frac{1}{4\sin^{2}\left(\frac{\pi(j+\theta)}{n}\right)-s}=\sum_{p\geq 0}s^{p}\sum_{j=0}^{n-1}\left[4\sin^{2}\left(\frac{\pi(j+\theta)}{n}\right)\right]^{-p-1}.$
@@
[Tr[$\mathcal{L}[\Tr[K_{n,\theta}]](s)=\sum_{p\geq 0}s^{p}\zeta_{n}(p+1,\theta)=\frac{1}{2}\frac{nU_{n-1}(1-s/2)}{T_{n}(1-s/2)-\cos(2\pi\theta)},$ (26)
@@
For any $0<\theta<1$ and for any integer $p\geq 1$, the map
$n\mapsto\zeta_{n}(p,\theta)$
is a polynomial function of $n$.
@@
It is a direct consequence of Lemma 6.4, since the generating function of $\zeta_{n}(p,\theta)$ satisfies
$\sum_{p\geq 0}s^{p}\zeta_{n}(p+1,\theta)=\frac{nU_{n-1}(1-s/2)}{2(T_{n}(1-s/2)-\cos(2\pi\theta))}.$
By the equation (25), both the numerator and the denominator are in $(\mathbb{C}[n])[[s]]$, with the denominator invertible in this ring since $T_{n}(1)=1$. ∎
@@ 6.4 Proof of the theorems and their corollaries
First recall from the asymptotic of Theorem 2.1 that when $s\in\mathbb{C}$ is not an half integer, $0<\theta<1$, $m\geq 0$, the following holds as $n\to\infty$;
$\zeta_{n}(s,\theta)=n\zeta_{\mathbb{Z}}(s)+n^{2s}\left(\sum_{k=0}^{m}a_{k}(s)\zeta_{\mathbb{R}/\mathbb{Z}}(s-k,\theta)n^{-2k}+\mathcal{O}(n^{-2-2m})\right).$ (27)
@@
Since $\zeta_{\mathbb{R}/\mathbb{Z}}$ vanishes at negative integers, one could expect that the asymptotic becomes an exact formula when $s$ is a positive integer. Indeed, this is the formula (23). It will be a direct consequence of Corollary 6.5.
@@
When $s=p$ is a positive integer, the left-hand side of the asymptotic of (27) is a polynomial in $n$. Since $m\geq 0$ can be chosen arbitrarily, this implies that the asymptotic expansion terminates and becomes an exact formula. As $\zeta_{\mathbb{Z}}$ vanishes at positive integers, see [FK17, Theorem 2.2], the formula (23) follows and therefore Theorem 1.2. ∎
@@
Fix $n,p\geq 1$ and let $0<\theta<1$. In the left hand side of formula (23), the term $j=0$ can be expanded as
$\left[4\sin^{2}\left(\frac{\pi\theta}{n}\right)\right]^{-p}=\sum_{k=0}^{p}a_{p-k}(p)\left(\frac{n}{2\pi\theta}\right)^{2k}+\mathcal{O}(\theta^{2}),$
@@
$\sum_{j=1}^{n-1}\left[4\sin^{2}\left(\frac{\pi(j+\theta)}{n}\right)\right]^{-p}$ $=\sum_{k=0}^{p}a_{p-k}(p)\zeta_{\mathbb{R}/\mathbb{Z}}(k,\theta)n^{2k}-\left[4\sin^{2}\left(\frac{\pi\theta}{n}\right)\right]^{-p}$
$=\sum_{k=0}^{p}a_{p-k}(p)(\zeta_{\mathbb{R}/\mathbb{Z}}(2k,\theta)-(2\pi\theta)^{-2k})n^{2k}+\mathcal{O}(\theta^{2}).$
Now one can take $\theta=0$ and the desired result follows. ∎
@@
One can either say that by the Corollary 6.5 the map $n\mapsto L_{n}(p,\chi)$ is a polynomial and therefore the asymptotic in Theorem 3.2 and in Theorem 1.1 must stops and becomes exact formulas. One can also directly sum the formulas in Theorem 1.2 against the character $\chi$ and by the standard formulas concerning $L$-functions, namely
and
$\sum_{r=1}^{q-1}\chi(r)(\partial_{\theta}^{a}\zeta_{\mathbb{R}/\mathbb{Z}})(s,r/q)=(-4s\pi)^{a}\cdot 2\left(\frac{q}{2\pi}\right)^{2s+a}L(2s+a,\chi),$
where $a=0$ if $\chi$ is even and $a=1$ if $\chi$ is odd. ∎

### v2

@@ 3 Asymptotics of discrete Dirichlet $L$-functions
$\zeta_{\mathbb{Z}/n\mathbb{Z}}(s)$ $=n\zeta_{\mathbb{Z}}(s)+n^{2s}\left(\sum_{k=0}^{l}a_{k}(s)\zeta_{\mathbb{R}/\mathbb{Z}}(s-k)n^{-2k}+\mathcal{O}(n^{-2-2l})\right).$
@@
The following asymptotic expression is obtained by Friedli in [F16].
@@
Let $\chi$ be an even, non-principal Dirichlet character of modulus $q>1$. Then, for any fixed $s\in\mathbb{C}\setminus\{\frac{1}{2},\frac{3}{2},\frac{5}{2},\ldots\}$ and any integer $m\geq 0$, the following asymptotic expansion holds as $n\to\infty$:
$L_{n}(s,\chi)=2\left(\frac{2\pi}{qn}\right)^{-2s}\left[\sum_{k=0}^{l}a_{k}(s)\left(\frac{2\pi}{qn}\right)^{2k}L(2(s-k),\chi)+\mathcal{O}(n^{-2-2l})\right],$
@@
$=n^{2s}\left(\sum_{k=0}^{l}a_{k}(s)\left(\frac{1}{-2\pi s}\right)^{a}\sum_{r=1}^{q-1}\chi(r)(\partial_{\theta}^{a}\zeta_{\mathbb{R}/\mathbb{Z}})(r/q,s-k)n^{a-2k}+\mathcal{O}(n^{a-2-2l})\right)$
$=2n^{2s}\left(\sum_{k=0}^{l}a_{k}(s)\left(\frac{(2s-k)}{s}\right)^{a}\left(\frac{q}{2\pi}\right)^{2(s-k)+a}L(a+2(s-k),\chi)n^{a-2k}+\mathcal{O}(n^{a-2-2l})\right)$
$=2\left(\frac{qn}{2\pi}\right)^{2s}\left(\sum_{k=0}^{l}a_{k}(s)\left(\frac{2(s-k)}{s}\right)^{a}L(a+2(s-k),\chi)\left(\frac{qn}{2\pi}\right)^{a-2k}+\mathcal{O}(n^{a-2-2l})\right).$
This concludes the proof of both Theorem 1.1 and Theorem 3.2, recalling that
@@
In fact, in the next section, what will be most important for us is the two first terms of the asymptotic, that is the case $l=1$ in Theorem 1.1. Precisely the asymptotic of size $l=1$ is
@@
Let $\chi$ be a non-principal, even and real Dirichlet character of modulus $q>1$. Fix $0<s<1$. Then $L(s-2,\chi)<0$ and it follows from the asymptotics in Theorem 3.2 that if $L_{n}(s,\chi)\geq 0$ for a subsequence of $n\rightarrow\infty$, then $L(s,\chi)>0$ and in particular non-zero.
Let $\chi$ be an odd and real Dirichlet character of modulus $q$. Fix $0<s<1$. The second term in the asymptotics in Theorem 1.1 now has a different sign, since (s3)/6<0(s-3)/6<0. It follows that if $L_{n}(s,\chi)\geq 0$ for a subsequence of $n\rightarrow\infty$, then $L(s,\chi)\geq 0$.
Since Siegel zeros are known to be simple for real characters, they cannot exist if $L(s,\chi)\geq 0$ for all $0<s<1$ (or actually, it is enough with a certain interval $1-\epsilon<s<1$.)
@@
and it satisfies the functional equation
@@
We now specialize the previous theorem when $\chi$ is real, odd and for real zeros of $L$. Recall that for such a $\chi$, Gauss himself proved that $w_{\chi}=1$. Therefore, as in the previous proof, for real $0<s<1$, if $\xi(s,\chi)\neq 0$, then
@@ 5.1 No real zero for Dirichlet $L$-functions with heat positive character
The Mellin transform of the sum in the hypothesis gives $L_{n}(s,\chi)\cdot\Gamma(s)$. The $\Gamma$ factor is positive on $s>0$, and hence heat positivity implies that $L_{n_{i}}(s,\chi)>0$. Therefore, since $L(s-2,\chi)<0$ for $0<s<1$ as remarked above, the asymptotics (12) force $L(s,\chi)>0$. ∎
@@ 5.2 First examples of heat positive characters
$(\Psi_{n}\rho)(\omega)=4\sin^{2}\left(\frac{\pi\omega}{qn}\right)\rho(\omega).$
One computes for $\omega\in\mathbb{Z}/qn\mathbb{Z}$,
@@
since we consider $\mathbb{Z}/qn\mathbb{Z}$ as discrete, the dual measure (with respect to Fourier) is therefore the compact one on $\mathbb{Z}/qn\mathbb{Z}$. Taking the trace of this solution viewed as an operator acting on $L^{2}(\mathbb{Z}/qn\mathbb{Z})$ by convolution amounts to putting $x=0$ and multiplying by $qn$, which gives
@@ 5.3 Further examples of heat positive characters
Consider the operator $S\coloneqq 3-\Delta$, defined by $(Sf)(x)\coloneqq f(x-1)+f(x)+f(x+1)$, which is the generator of the random walk on $\mathbb{Z}/N\mathbb{Z}$ allowing the walk to stand still. Let us also consider $\partial$, the forward difference operator, defined by $(\partial f)(x)\coloneqq f(x+1)-f(x)$. Then $K(t)=e^{-3t}\exp(tS)\delta_{0}$, and therefore we need only to analyze $\phi(t)\coloneqq\exp(tS)\delta_{0}$. But clearly,
@@ 6 Special values
In this section, we prove Theorem 1.2, and by doing so we link special values of both discrete zeta functions and discrete $L$-functions with special values of the Riemann zeta function and Dirichlet $L$-functions, respectively. For this, we first use the Laplace transform of the discrete heat kernel on the $\theta$-bundle over $\mathbb{Z}/n\mathbb{Z}$ together with a precise expression of the characteristic polynomial of the bundle Laplacian $\Delta_{n,\theta}$ (namely, the Chebyshev polynomials) to obtain a polynomial behavior of the discrete special value. Combining this polynomial behavior with the asymptotics of Theorem 2.1, we obtain closed formulas for the zeta functions associated. By using standard formulas for $L$-functions and specializing $\theta$ accordingly, we obtain the desired relation between discrete and continuous zeta and $L$-functions.
@@ 6.1 Formulas
In terms of discrete zeta function on the $\theta$-bundle, Theorem 1.2 says that for any $0<\theta<1$ and integers $n,m\geq 1$, we have
$\zeta_{\mathbb{Z}/n\mathbb{Z}}(m,\theta)=\sum_{k=0}^{m}a_{m-k}(m)\zeta_{\mathbb{R}/\mathbb{Z}}(k,\theta)n^{2k}.$ (25)
As above, the coefficients $m\mapsto a_{k}(m)$ are polynomial functions of degree $k$, given by the Taylor expansion
@@
For any integers $n,m\geq 1$, the following holds:
$\zeta_{\mathbb{Z}/n\mathbb{Z}}(m)=\sum_{k=0}^{m}a_{m-k}(m)\zeta_{\mathbb{R}/\mathbb{Z}}(k)n^{2k}.$
@@ 6.3 Laplace transform of the bundle heat kernel
$\sum_{j=0}^{n-1}\frac{1}{4\sin^{2}\left(\frac{\pi(j+\theta)}{n}\right)-s}=\sum_{m\geq 0}s^{m}\sum_{j=0}^{n-1}\left[4\sin^{2}\left(\frac{\pi(j+\theta)}{n}\right)\right]^{-m-1}.$
@@
[Tr[$\mathcal{L}[\Tr[K_{n,\theta}]](s)=\sum_{m\geq 0}s^{m}\zeta_{n}(m+1,\theta)=\frac{1}{2}\frac{nU_{n-1}(1-s/2)}{T_{n}(1-s/2)-\cos(2\pi\theta)},$ (28)
@@
As a result, we obtain the polynomial behavior of the special value of $\zeta_{\mathbb{Z}/n\mathbb{Z}}(m,\theta)$.
@@
For any $0<\theta<1$ and for any integer $m\in\mathbb{Z}$, the map
$n\mapsto\zeta_{n}(m,\theta)$
is a polynomial function on the domain $\{n\in\mathbb{N}\mid n>-m\}$.
@@
The case of positive $m\geq 1$ is a direct consequence of Lemma 6.4. Indeed, for each $n\geq 1$ the generating function of the numbers $\zeta_{n}(m,\theta)$, $m\geq 1$ converges for some $s\neq 0$ and satisfies
$\sum_{m\geq 1}s^{m}\zeta_{n}(m,\theta)=s\cdot\frac{nU_{n-1}(1-s/2)}{2(T_{n}(1-s/2)-\cos(2\pi\theta))}.$
By the equation (27), both the numerator and the denominator of the right hand side are in $(\mathbb{C}[n])[[s]]$, with the denominator invertible in this ring since $T_{n}(1)=1$.
For non-positive $m\leq 0$, let $n>m$ and expand the sine function into exponential
$\zeta_{n}(m,\theta)$ $=4^{m}\sum_{j=0}^{n-1}\left[\sin\left(\frac{\pi(j+\theta)}{n}\right)\right]^{2m}$
$=\sum_{k=-r}^{r}\binom{2m}{m+k}(-1)^{k}e^{2\pi\theta k/n}\sum_{j=0}^{n-1}e^{2\pi jk/n}$
$=n\binom{2m}{m}$.
Observe that in the last calculation, we have explicitly shown that for $m\leq 0$,
$\zeta_{\mathbb{Z}/n\mathbb{Z}}(-m,\theta)=n\zeta_{\mathbb{Z}}(-m),$
for all $0<\theta<1$, and all $n>m$.
@@ 6.4 Proof of the theorems and their corollaries
First recall from the asymptotic of Theorem 2.1 that when $s\in\mathbb{C}$ is not a half integer, $0<\theta<1$, $l\geq 0$, the following holds as $n\to\infty$;
$\zeta_{n}(s,\theta)=n\zeta_{\mathbb{Z}}(s)+n^{2s}\left(\sum_{k=0}^{l}a_{k}(s)\zeta_{\mathbb{R}/\mathbb{Z}}(s-k,\theta)n^{-2k}+\mathcal{O}(n^{-2-2l})\right).$ (29)
@@
Since $\zeta_{\mathbb{R}/\mathbb{Z}}$ vanishes at negative integers, one could expect that the asymptotic becomes an exact formula when $s$ is an integer. In the case of positive $s=m\geq 1$ this is the formula (25). It will be a direct consequence of Corollary 6.5, and in fact, as already proven, this phenomenon holds as well for non-positive $m\leq 0$ and takes the form (4).
@@
When $s=m$ is a positive integer, the left-hand side of the asymptotic of (29) is a polynomial in $n$. Since the length $l\geq 0$ can be chosen arbitrarily, this implies that the asymptotic expansion terminates and becomes an exact formula. As $\zeta_{\mathbb{Z}}$ vanishes at positive integers, as $\zeta_{\mathbb{Z}}$ is the analytical continuation of the binomial coefficients, the formula (25) follows and therefore Theorem 1.2. ∎
@@
Fix $n,m\geq 1$ and let $0<\theta<1$. In the left hand side of formula (25), the term $j=0$ can be expanded as
$\left[4\sin^{2}\left(\frac{\pi\theta}{n}\right)\right]^{-m}=\sum_{k=0}^{m}a_{m-k}(m)\left(\frac{n}{2\pi\theta}\right)^{2k}+\mathcal{O}(\theta^{2}),$
@@
$\sum_{j=1}^{n-1}\left[4\sin^{2}\left(\frac{\pi(j+\theta)}{n}\right)\right]^{-m}$ $=\sum_{k=0}^{m}a_{m-k}(m)\zeta_{\mathbb{R}/\mathbb{Z}}(k,\theta)n^{2k}-\left[4\sin^{2}\left(\frac{\pi\theta}{n}\right)\right]^{-m}$
$=\sum_{k=0}^{m}a_{m-k}(m)(\zeta_{\mathbb{R}/\mathbb{Z}}(2k,\theta)-(2\pi\theta)^{-2k})n^{2k}+\mathcal{O}(\theta^{2}).$
Now one can take the limit $\theta\to 0$ and the desired result follows. ∎
@@
Since
$\sum_{r=1}^{q-1}\chi(r)(\partial_{\theta}^{a}\zeta_{n})(s,r/q)=\left(-s\frac{2\pi}{n}\right)^{a}\cdot L_{n}(s,\chi),$
for $a\in\{0,1\}$ being the same parity as $\chi$, the Corollary 6.5 implies that for each $m\geq 1$ the map $n\mapsto L_{n}(m,\chi)$ is a polynomial and therefore the asymptotic in Theorem 3.2 and in Theorem 1.1 must stops and becomes exact formulas. For $m=0$ and $\chi$ even, the corresponding formula of Corollary 1.4 is trivial since both side are 0. For $m=0$ and $\chi$ odd, one has
$L_{n}(0,\chi)$ $=\lim_{s\to 0}\frac{n}{2\pi}\frac{1}{s}\sum_{r=1}^{q-1}\chi(r)(\partial_{\theta}\zeta_{n})(s,r/q)$
$=\frac{n}{2\pi}\sum_{r=1}^{q-1}\chi(r)\frac{\partial}{\partial\theta}\Big|_{\theta=r/q}\zeta_{n}^{\prime}(0,\theta).$
But the derivative at 0 of the spectral zeta function associated to the $\theta$-bundle is the log-determinant of $\Delta_{n,\theta}$. That is
$\zeta_{n}^{\prime}(0,\theta)=-\log\det(\Delta_{n,\theta})=-\log(4\sin^{2}(\pi\theta)).$
In particular, it does not depend on $n$. Thus, $n\mapsto L_{n}(0,\chi)$ is a polynomial and the same argument follows. One could have argue alternatively that
$L_{n}(0,\chi)$ $=\sum_{j=1}^{qn-1}\chi(j)\cot\left(\frac{\pi j}{qn}\right)$
$=\sum_{r=1}^{q-1}\chi(r)\sum_{j=0}^{n-1}\cot\left(\frac{\pi r}{qn}+\frac{\pi j}{n}\right)$
$=n\sum_{r=1}^{q-1}\chi(r)\cot\left(\frac{\pi r}{q}\right),$
using a well-known cotangent formula. This results in the same polynomial behavior. ∎